solve for x: 35-(3-22x)=2[10x-3(x+4)]
Answer by oscargut(2103)

solve for x:

35-(3-22x)=2[10x-3(x+4)]
35-3+22x = 2(10x-3x-12)
22x+32 = 14x-24
8x = -56
Answer: x = -7